• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Peter L

    Log entry time 00:26:30 on July21,2005

    Entry number 148814

    keyword=Swing Summary

    We started with the right septum down and began to recover.  David A and JP 
    went down into the hall to change the target density, and David L. needed to do 
    Compton work.
    We took a harp scan.
    
    We then changed courses and switched to the Spec DAQ and set the BPMA to 1.6, 
    0.0 and BPMB to -1.4, 0.0  The rates in the two Happex detectors settled within 
    5% of eachother with the He target.
    
    Next we changed the momentum setting to -2% and ran one nominal and one run 
    with the s2's turned off.
    2181 -2% Po He S2 on
    2182 -2% Po He S2 off
    2183 -7% Po Carbon target
    2184 -7% 21cm
    2185 -7% Al dummy
    2186 -7% carbon foil
    2187-2190 Profile scans
    2191 ps=1